[figure: 9.7 cm]

COERCIVE PROPERTIES OF Co/Gd MULTILAYERS WITH Si

SPACER

A.B. Chizhik^{1}, S.L. Gnatchenko^{1}, D.N. Merenkov^{1}, V.O. Vas’kovskiy^{2}, A.V. Svalov^{2}

1 Institute for Low Temperature Physics & Engineering, 310164, Kharkov, Ukraine

2 Institute of Physics & Applied Mathematics, Ural State University, 620083, Ekaterinburg, Russia

Investigation of magnetization reversal of rf sputtered (Co30Å/SiXÅ/Gd75Å)·20 multilayers (X = 0, 2, 3, 5, 10 Å) has been performed in the temperature range 20-300K using the longitudinal magneto-optical Kerr effect. Due to the ferrimagnetic ordering of magnetic moments of Co and Gd layers, the compensation phenomenon has been observed. The sign of rotation of the light polarization plane was different for T<T_{C} and for T>T_{C} (T_{C} is compensation temperature), where the multilayers were Gd-aligned with the field or Co-aligned, respectively. The value of T_{C} depended upon Si spacer thickness X. It decreased from T_{C}=330K (X=0) to T_{C}=138K (X=5Å), disappearing for X=10Å. The change of the compensation temperature is associated with the change of the influence of Co subsystem on Gd subsystem. When approaching T_{C}, an increase of coercivity has been observed. This effect is originated by the decrease of the net magnetization near the compensation point. The investigation demonstrated the possibility to change the value of T_{C} and coercivity of multilayers, by changing the thickness of the nonmagnetic spacer.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : A.B. Chizhik

Address for correspondence : Institute for Low Temperature Physics & Engineering, 47 Lenin Ave., 310164 Kharkov, UKRAINE

Email address : hizhik@ilt.kharkov.ua